OpenStudy (anonymous):

Please Help Me!!! If f(x) = 2x2 - 30, find f(4). 35 -14 2 17

OpenStudy (texaschic101):

f(x) = 2x^2 - 30 f(4) = 2(4^2) - 30 f(4) = 2(16) - 30 f(4) = 32 - 30 f(4) = 2 basically, your just subbing in 4 for x

OpenStudy (anonymous):

oh, thats what i forgot to do, i was having trouble because i forgot to substitute for BOTH x's thanx so much

OpenStudy (texaschic101):

no problem :)